OpenStudy (anonymous):

4m+n=35 m-5n=14... use substitution method to solve.

OpenStudy (anonymous):

okay: (1) 4m+n=35 (2) m-5n=14 -> m = 14+5n sun (2) into (1) 4(14+5n) + n = 35 56 + 20n +n = 35 21n = -21 n = -1 sub n = -1 back into (2) m = 14 + 5(-1) m = 14 -5 m = 9

OpenStudy (anonymous):

Thank you so much Ibd05 I am going to try to see if I understand this, if I don;t is it okay if I ask you some questions to see how you solved this?

OpenStudy (anonymous):

yup, ask away :)